Supppose that a farm's production function is determined by two factors: land area (K) and labour (L). The
production function is specified as follows:
Q = 64K^0.5 L^0.5
The farm's land area is 500 acres
The unit cost of capital and labour is as follows:
- land $10 per acre
- labour $500 per labour

What is Q when L is 1,2,3,4 and 5?
1
Expert's answer
2017-11-24T13:57:06-0500
Q = 64K^0.5 L^0.5,
S = 500 acres,
k = $10 per acre,
w = $500 per labour.
If L = 1, then Q = 64*500^0.5*1^0.5 = 1431.08 units,
If L = 2, then Q = 64*500^0.5*2^0.5 = 2023.86 units,
If L = 3, then Q = 64*500^0.5*3^0.5 = 2478.71 units,
If L = 4, then Q = 64*500^0.5*4^0.5 = 2862.17 units,
If L = 5, then Q = 64*500^0.5*5^0.5 = 3200 units.